In doing a five-year analysis of future dividends, the Dawson Corporation is considering the following two plans. The values represent dividends per share. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

  

  Year Plan A Plan B
  1 $ 1.10 $ .10
  2 1.10 1.60
  3 1.10 .20
  4 1.50 3.50
  5 1.50 1.60
a.

How much in total dividends per share will be paid under each plan over five years? (Do not round intermediate calculations and round your answers to 2 decimal places.)

  

Total Dividends
  Plan A $   
  Plan B $   

  

b-1.

Mr. Bright, the Vice-President of Finance, suggests that stockholders often prefer a stable dividend policy to a highly variable one. He will assume that stockholders apply a lower discount rate to dividends that are stable. The discount rate to be used for Plan A is 11 percent; the discount rate for Plan B is 13 percent. Compute the present value of future dividends. (Do not round intermediate calculations and round your answers to 2 decimal places.)

  

Present Value of
Future Dividends
  Plan A $
  Plan B $
Answer #1

a) Total dividends payable for both plans are:

Plan A: $1.1 + $1.1 + $1.1 + $1.5 + $1.5 = $6.3

Plan B: 0.10 + 1.60 + 0.20 + 3.50 + 1.60 = $7

b) PV of dividends in both plans. I am calculating dividends for Part A using the formula methods and Part B using the financial calculator method. This would help you understand both the approaches (please practice the other approaches for both plan).

Plan A (formula approach):

PV = \frac{1.10}{(1 + 0.11)^1} + \frac{1.10}{(1 + 0.11)^2} + \frac{1.10}{(1 + 0.11)^3} + \frac{1.5}{(1 + 0.11)^4} + \frac{1.50}{(1 + 0.11)^5}

PV = 0.991 + 0.893 + 0.804 + 0.988 + 0.890

PV = $4.57

Plan B (calculator approach):

We would use the CF - NPV functions for calculating the present value of stream of dividends. Below are the steps:

Step 1: It is agood practice to always begin by clearing the Cash Flow worksheet. Press [CF] to turn the CF worksheet on. Press [2nd] [CLR] [Quit]

Step 2: Press [CF]. The display should show: CF0= 0.00000. Now, this is zero in our question, so leave it.

Step 3: Press the down arrow key to display C01. Type in the cash amount for period 1 which is 0.10, then press [Enter]. Press the down arrow key to display F01. Press [Enter] to accept the default amount of 1.00000.

Step 4: Press the down arrow key to display C02. Type in the cash amount for period 2 which is 1,60, then press [Enter]. Press the down arrow key to display F02. Press the down arrow key again to accept F02, and to move to the next field which will display as C03.

Step 6: Type in the cash amount for each subsequent period in the same manner.

Step 7: To compute NPV: Press [NPV] to display I = 0.0000. Enter the required rate of return 13. Press [Enter]. Press the down arrow key, then press [CPT] to display the dollar amount of the NPV. Answer displayed is $4.49517

Hence, answer is $4.50

Hope these steps are easy to understand. Let me know if not.